# Spin polarized enantio-sensitive multipolar photoelectron currents

### v3

@@ Abstract
This indicates that $\boldsymbol{\vec{j}}_{\tau}^{L}$ belongs to the Class III enantio-sensitive observables. Hence, to find spin-sensitive effects, we have to consider multipolar photoelectron currents, i.e. currents resolved along two orthogonal directions. Without spin detection, multipolar enantio-sensitive currents could be induced using orthogonally polarised two-color fields [13, 33, 26, 5], and recorded via interference between the two photoionization pathways involving the interference of one-photon and two-photon transitions.
The three components of $\boldsymbol{\vec{j}}_{\text{circ}}^{L}$ become mutually orthogonal when $\boldsymbol{\hat{s}}\perp\boldsymbol{\hat{\Xi}}$, hence, can be measured independently [see Fig. 2(d-f)]. A comparison of each term in $\boldsymbol{\vec{j}}_{\text{circ}}^{L}$ is shown in Fig. 4. The vortex current $\boldsymbol{\vec{j}}_{\tau}^{L}$ is negligible for the state $|\psi_{1,\frac{1}{2}}^{\pm}\rangle_{c}$ while the radial current $\boldsymbol{\vec{j}}_{r}^{L}$ is almost eight times than PECD. Meanwhile, $\boldsymbol{\vec{j}}_{r}^{L}$ is negligible for the state $|\psi_{-1,\frac{1}{2}}^{\pm}\rangle_{p}$ but the vortex current $\boldsymbol{\vec{j}}_{\tau}^{L}$ is almost twice the PECD. Thus, spin detection opens an exciting opportunity to detect enantio-sensitive multipolar currents already in the one-photon regime, because the orthogonal axes can be marked by the photoelectron momentum and its spin.
In summary, we have shown how measurements of the photo-electron spin can be used to detect spin polarized enantio-sensitive currents. These currents are enabled by two geometric mechanism: (i) the Bloch pseudovector $\boldsymbol{\vec{\mathcal{S}}}_{\boldsymbol{\vec{k}}}$ which enables collinear locking of the photoelectron current with the photoelectron spin and operators for arbitrary light fields, even isotropic illumination, and (ii) the spin-resolved propensity field $\vec{\mathbb{B}}_{\boldsymbol{\vec{k}}}$ which enables orthogonal locking and only operators for light fields with a fixed polarization plane. We have also presented several detection schemes to measure these currents.

### v4

@@ Abstract
This indicates that $\boldsymbol{\vec{j}}_{\tau}^{L}$ belongs to the Class III enantio-sensitive observables predicted in Ref. [23]. Without spin detection, multipolar enantio-sensitive currents could be induced using orthogonally polarised two-color fields [11, 31, 26, 3], and recorded via interference between the two photoionization pathways involving the interference of one-photon and two-photon transitions.
The three currents contributing to $\boldsymbol{\vec{j}}_{\text{circ}}^{L}$ (Eq.16) become mutually orthogonal when $\boldsymbol{\hat{s}}\perp\boldsymbol{\hat{\Xi}}$. In this case they can be measured independently (see Fig. 3). A comparison of each term in $\boldsymbol{\vec{j}}_{\text{circ}}^{L}$ is shown in Fig. 4c. The vortex current $\boldsymbol{\vec{j}}_{\tau}^{L}$ is negligible for the state $|\psi_{1,\frac{1}{2}}^{\pm}\rangle_{c}$ while the radial current $\boldsymbol{\vec{j}}_{r}^{L}$ is almost eight times higher than PECD. Meanwhile, $\boldsymbol{\vec{j}}_{r}^{L}$ is negligible for the state $|\psi_{-1,\frac{1}{2}}^{\pm}\rangle_{p}$ but the vortex current $\boldsymbol{\vec{j}}_{\tau}^{L}$ is almost twice as large as the PECD.
We have established two fundamental mechanisms of CISS in photoionization operating in time-even and time-odd (i.e. involving photon spin) set-ups: (i) the momentum -resolved Bloch pseudovector $\boldsymbol{\vec{\mathcal{S}}}_{\boldsymbol{\vec{k}}}$ which enables collinear locking of the photoelectron current with the photoelectron spin and operates in arbitrary light fields, even under isotropic illumination, and (ii) the spin-resolved propensity field $\vec{\mathbb{B}}_{\boldsymbol{\vec{k}}}$ which enables orthogonal triple locking between the photoelectron spin, its current and the photon spin.
We find [20] that the spin-conditioned photoelectron currents derived in Eqs. (13a), (17), and (20a) are consistent with the kinematic predictions of Cherepkov [7, 8, 10]. Our treatment makes explicit the dynamical mechanisms that were implicit in those earlier works. Importantly, it also leads to a qualitatively new prediction: the generation of a spin conditioned current even for randomly oriented chiral molecules subjected to isotropic illumination.
O.S., A.F. O. and P.C.F. acknowledge ERC-2021-AdG project ULISSES, grant agreement No 101054696. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. A.F.O. acknowledges funding from the Royal Society URF/R1/201333, URF/ERE/210358, and URF/ERE/231177 and from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - 543760364.